OpenStudy (anonymous):

Solve: 7(x − 1) = 5(x + 2) + 1 9 18 2 4

thomaster (thomaster):

First get rid of the parenthesis, can you do that?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ONLYLN 1D IM SORRY

OpenStudy (anonymous):

distributive property right

OpenStudy (anonymous):

YEA

OpenStudy (anonymous):

hmm ..

OpenStudy (anonymous):

sure you are

thomaster (thomaster):

@ASVPJOHNNY please don't spoil the post

OpenStudy (anonymous):

IM SEROUS MY GIRL TOLD ME TO SAY SORRY SO I DID

OpenStudy (anonymous):

thats what i got down on my paper

OpenStudy (anonymous):

ONLYLN 1D DO U FOREGIVE ME

mathslover (mathslover):

Shkrina, please let thomaster discuss the quest. with Onlyin1d . Interrupting in between may cause both the answerer and the asker distract from the right way.

thomaster (thomaster):

Yea 7x-7=5x+10+1 do you know what do do next?

OpenStudy (anonymous):

@mathslover sure ...

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

PLEZZZZ

thomaster (thomaster):

First you combine x's and the numbers You can't currently combine any x's. Where can you combine numbers?

OpenStudy (anonymous):

no you. im trying to work so can you plz not disturb with your comments.

OpenStudy (anonymous):

WHERE CAN U COMBINE NUMBERS

OpenStudy (anonymous):

12x = 3+1

OpenStudy (anonymous):

OK SO WHAT DOES THAT GIVE U

thomaster (thomaster):

\(\Large\sf 7x-7=5x+10+1\) you can combine the 10 and the 1 to get: \(\Large\sf 7x-7=5x+11\) Now you bring the numbers to the right and the x's to the left

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

IT IS IM SORRY PLEZZZ

thomaster (thomaster):

No it won't help to just give the answer, you need to learn how to do it yourself. Do you know how to bring x's to the left? \(\Large\sf 7x-7=5x+11\)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

you can do it

OpenStudy (anonymous):

ok now what

OpenStudy (anonymous):

U HAVE TO DOAT HE GAVE U

thomaster (thomaster):

Now you bring x's to the left and numbers to right

OpenStudy (anonymous):

is it 18

OpenStudy (anonymous):